Find the general solution of the differential equation: y'−4y=te^−2t

Use lower case c for the constant in your answer.
Answer #1

Ans:- Given -25. y-ay=té2t 2t. => dy la ay = teht eladta ett I.FC => dat lyett) Aty=teht. ett l-442=te? • Stéstst=ts 0. now,36 From , + ren differed heral solution of the 6t. (6t+lj y=-36 (6E+1) where e=clett is is arbitra constant differedntial equ